OpenStudy (anonymous):

what is the equation of the line passing through (-3,6) and parrellel to the line y=4x-2 in standard form?

jimthompson5910 (jim_thompson5910):

Anything parallel to y = 4x-2 has a slope of 4 y = mx+b y = 4x+b 6 = 4(-3)+b 6 = -12+b 6+12=b 18 = b b = 18 So the equation of the line is y = 4x+18 Now convert to standard form y = 4x+18 -4x+y = 18 4x-y = -18